Evaluate the following definite integral:

\int_{\frac{1}{2}}^{1}\frac{e^{-\frac{1}{y^2}}}{y^3}dy
Answer #1

Soly e yn dy 93 simit when y=1/ to those det let y2 diff.w.rtott edt It=-4 when y=1 -C-243) dy=de dya Idy = yedt dt t=-* 2 y3